# On the Importance of Embedding Norms in Self-Supervised Learning

### v1

@@ 6 The Embedding-Norm Effect in Practice
To understand how the embedding-norm effect influences cosine-similarity-based SSL training, we investigate three distinct interventions which should mitigate it. These interventions provide controlled settings to analyze the empirical relationship between embedding norms and SSL training.
@@
Our first intervention mechanism—weight decay—is already present in essentially all SSL models. The idea here is that adding a penalty on the weights implicitly regularizes the embedding norms (Wang et al., 2017). Figure 5 demonstrates this effect in SimCLR and SimSiam training: without weight decay ($\gamma=0$), embeddings grow unconstrained, while excessive weight decay ($\gamma=5e\text{-}2$) causes collapse. With appropriate values ($\gamma=1e\text{-}5$ for SimCLR, $\gamma=5e\text{-}4$ for SimSiam), norms decrease gradually, leading to improved $k$-nn accuracy. Interestingly, the embedding norm’s impact on the $k$-nn accuracy is much more pronounced in the attraction-only setting. We also see that, even without weight decay, the embeddings can shrink (as occurs for SimSiam). We attribute this to the embeddings being produced by shared weights: although the gradients require all embeddings to grow, a shared weights matrix may not be able to independently update every point’s position.
@@
Figure 5: Effect of the weight decay on the embedding norms in SimCLR and SimSiam. Epochs are log-scale and go from 1 to 256. Corresponding $k$-nn accuracies are plotted in the bottom row.
Cut-Initialization.
As shown in Figure 5, weight decay gradually reduces embedding norms over time, but doesn’t control them at the start of training. To address this, we propose cut-initialization - dividing all network weights by a constant $c$ at initialization. This ensures small embedding norms at initialization which are then kept small via weight decay. We implement this uniformly across all layers for simplicity (Listing LABEL:alg:cut_init). Interestingly, a variant of this can be found in HuggingFace’s default image transformer code (PyTorch Foundation, 2025) and we know at least one cos.sim.-based SSL model which uses it (Peng et al., 2022).
@@
We study the interplay between cut-initialization and weight decay values on SimCLR and SimSiam in Table 1. Specifically, we report the $k$-nn classification accuracy after 100 epochs on the CIFAR-100 and ImageNet-100 datasets and see that, in both the contrastive and non-contrastive settings, pairing cut-initialization with weight decay accelerates the training process. As was the case for the weight decay, the difference is more stark in the non-contrastive setting. $c=2$ performed best for SimCLR, providing an additional 2% in accuracy at the default weight decay, while $c=8$ led to about a 10% increase for SimSiam. For the remaining experiments, we use $c=3$ for SimCLR and $c=9$ for SimSiam. We report a variant of Figure 5 including BYOL experiments in Figure S4. We furthermore show the $k$-nn classifier accuracies at $500$ epochs with and without cut-initialization in Table 2. Here we see that pairing SSL models with cut-initialization often helps the model reach higher final accuracies.
We also evaluate cut-initialization in imbalanced data settings. For Cifar-10, we use the exponential split from (Van Assel & Balestriero, 2024), where class $i$ has $n_{i}=5000\cdot 1.5^{-i}$ samples. Similarly for Cifar-100, the $i$-th class receives $n_{i}=500\cdot 1.034^{-i}$ samples. This way, all classes are represented and both imbalanced datasets contain roughly 15K samples. We also use Flowers102’s naturally long-tailed test set for training (Nilsback & Zisserman, 2008), evaluating on its validation set. Table 3 then shows that, in class-imbalanced settings, pairing SSL training with interventions on the embedding-norm effect can provide double-digit accuracy improvements.
Table 1: $k$-nn accuracy at epoch 100 for various values of cut-constant $c$ and weight decay $\lambda$. Left: SimCLR on Cifar-100. Right: SimSiam on ImageNet-100. Default weight-decay is underlined.
@@
GradScale Layer.
Perhaps the cleanest way to overcome the embedding-norm effect is to simply rescale the gradient directly. We achieve this using a custom PyTorch autograd.Function which we refer to as GradScale (for a full implementation, see Listing LABEL:alg:grad_scaling in the Appendix). This layer accepts a power parameter $p$ and is simply the identity function in the forward pass. However, the backwards pass multiplies each sample $z_{i}$’s contribution to the gradient by $\|z_{i}\|^{p}$. Thus, choosing power $p=0$ gives the default setting while choosing $p=1$ will cancel the gradients’ dependence on the embedding norm. We visualize the resulting gradient fields with powers $p=0$ and $p=1$ for a 2D embedding space in Figure 6. We refer to training with GradScale power $p=1$ simply as GradScale.
@@
Table 2 demonstrates that, when training remains stable, SimCLR’s $k$-nn accuracy benefits from GradScale’s embedding norm cancellation. Consistent with our cut-initialization experiments, this improvement becomes particularly pronounced on the class-imbalanced datasets in Table 3: GradScale provides a roughly 5% accuracy increase across the imbalanced datasets. While these results are promising, we observed that GradScale with $p=1$ failed to converge on Imagenet-100 and for non-contrastive models. This limitation aligns with these models’ known sensitivity issues (Zhang et al., 2022; Richemond et al., 2023).
Figure 6: Left: the default gradient field of the cosine similarity with respect to the north-pointing line (blue). Right: the gradient field using the GradScale layer with $p=1$.
Table 2: $k$-nn accuracies at epoch 500 for default, cut-initialized and GradScale training on standard image datasets.
Cifar-10 Cifar-100 Imagenet-100
SimCLR Default 85.2 51.9 59.4
Cut ($c=3$) 87.0 52.6 60.9
GradScale 86.5 54.0 01.0
SimSiam Default 87.0 61.1 62.0
Cut ($c=9$) 89.0 61.5 67.2
Table 3: $k$-nn accuracies at epoch 500 for default, cut-initialized and GradScale training on class-imbalanced image datasets.
@@
Takeaways.
These results make it clear that the embedding norm effect impacts SSL training – particularly in non-contrastive settings – and can be mitigated using our proposed strategies. The effect appears most detrimental in class-imbalanced settings, aligning with our results on SSL confidence: imbalanced data creates variance in embedding norms, destabilizing training. Nonetheless, there remain questions which are beyond the scope of this work:
@@
In addition to seeing that the embedding-norm effect is more pronounced in attraction-only settings, we have found that the embeddings can shrink even in the absence of weight decay. We attribute both phenomena to the network’s shared weights: while our theory predicts uniform embedding growth, producing these embeddings via a single set of weights creates competition between different regions of the latent space. This competition would be especially relevant for the InfoNCE loss, which enforces uniformity over the latent hypersphere (Wang & Isola, 2020).
@@
We have been careful to not describe the embedding-norm effect as a strictly negative phenomenon. Consider the common transfer-learning setting in which prototypical class examples should anchor the learned representation (Pan et al., 2019; Lee et al., 2022). Our findings suggest the embedding-norm effect naturally supports this goal: prototypical examples should have large embedding norms and consequently would receive smaller gradients.
@@ 7 Discussion
Given that these properties are natively present in widely-used SSL approaches, they suggest several research directions beyond those already stated in the paper. First, embedding norms could serve as simple yet effective reliability metrics during inference. Second, the embedding norms provide a new lens for studying the modality gap in multi-modal representation learning (Liang et al., 2022). Lastly, the interplay between the embedding norm’s roles–as both a confidence metric and gradient scalar–suggests training pipelines that explicitly leverage both mechanisms.
@@ Acknowledgements
Andrew Draganov is partially supported by the Independent Research Fund Denmark (DFF) under a Sapere Aude Research Leader grant No 1051-00106B. Sharvaree Vadgama is supported by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research (NWO). This work was partially funded by the Gemeinnützige Hertie-Stiftung, the Cyber Valley Research Fund (D.30.28739), and the National Institutes of Health (UM1MH130981). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. Dmitry Kobak is a member of the Germany’s Excellence cluster 2064 “Machine Learning — New Perspectives for Science” (EXC 390727645). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Jan Niklas Böhm.
@@ A.2 InfoNCE Gradients
$\nabla_{i}^{\mathcal{R}}$ $=-\frac{\sum_{k\not\sim i}\text{ExpSim}(z_{i},z_{k})\frac{\partial\frac{z_{i}^{\top}z_{k}}{\|z_{i}\|\cdot\|z_{k}\|}}{\partial z_{i}}}{\sum_{k\not\sim i}\text{ExpSim}(z_{i},z_{k})}$
$=-\frac{\sum_{k\not\sim i}\text{ExpSim}(z_{i},z_{k})\frac{\partial\frac{z_{i}^{\top}z_{k}}{\|z_{i}\|\cdot\|z_{k}\|}}{\partial z_{i}}}{S_{i}}$
@@
We sum this this with the gradients of the attractive term to obtain the full InfoNCE gradient, completing the proof. ∎
We note that the repulsive force is weighted average over a set of unit vectors. Consequently, the repulsive gradient is smaller than the attractive one. Additionally, we point out that these gradients are symmetric: just like positive and negative samples $z_{j}$ and $z_{k}$ affect $z_{i}$, $z_{i}$ affects $z_{j}$ and $z_{k}$.
@@
First, consider that we applied the cosine similarity’s gradients from Proposition 1. Since $z_{i}$ and $(z_{j})_{\perp z_{i}}$ are orthogonal, $\|z_{i}^{\prime}\|_{2}^{2}=\|z_{i}\|^{2}+\frac{\gamma^{2}}{\|z_{i}\|^{2}}\|(z_{j})_{\perp z_{i}}\|^{2}$. The second term is positive if $\sin\phi_{ij}>0$.
@@ B.1 Aparametric Simulations
In essence, when $\alpha=0$, $x_{2}^{\prime}=x_{2}$. However, when $\alpha=1$ (resp. $\alpha=-1$), $x_{2}^{\prime}=x_{1}$ (resp. $x_{2}^{\prime}=-x_{1}$).
@@ B.2 Parametric Simulations
We restate the entire implementation for the simulations in Section 4.2 for completeness. We choose centers for 4 latent classes $\{c_{1},c_{2},c_{3},c_{4}\}$ uniformly at random from $\mathbb{S}^{10}$ by randomly sampling vectors from a standard multivariate normal distribution and normalizing them to the hypersphere. We then obtain the latent samples $\tilde{z}$ around center $c_{i}$ via $z\sim\mathcal{N}(c_{i},0.1\cdot\mathbf{I})$ and re-normalizing to the hypersphere. For each center, we produce 1K latent samples; these constitute our latent classes. We depict an example of 8 such latent classes (in 3 dimensions) in Figure 2(a). We finally obtain the dataset by generating a random matrix in $\mathbb{R}^{11\times 64}$ and applying it to the latent samples.

### v2

@@ 6 The Embedding-Norm Effect in Practice
To understand how the embedding-norm effect influences cosine-similarity-based SSL training, we investigate three distinct interventions which should mitigate it. These interventions provide controlled settings to analyze the empirical relationship between embedding norms and SSL training. Our experiments are on the Cifar-10, Cifar-100, Imagenet-100 and Tiny-Imagenet (Le & Yang, 2015) datasets.
@@
Our first intervention mechanism — weight decay — is already present in essentially all SSL models. The idea here is that adding a penalty on the weights implicitly regularizes the embedding norms (Wang et al., 2017). Figure 5 demonstrates this effect in SimCLR and SimSiam training: without weight decay ($\gamma=0$), embeddings grow unconstrained, while excessive weight decay ($\gamma=5\cdot 10^{-2}$) causes collapse. With appropriate values ($\gamma=10^{-5}$ for SimCLR, $\gamma=5\cdot 10^{-4}$ for SimSiam), norms decrease gradually, leading to improved $k$NN accuracy. Interestingly, the embedding norm’s impact on the $k$NN accuracy is much more pronounced in the attraction-only setting. We also see that, even without weight decay, the embeddings can shrink (as occurs for SimSiam). We attribute this to the embeddings being produced by shared weights: although the gradients require all embeddings to grow, a shared weights matrix may not be able to independently update every point’s position.
@@
Figure 5: Effect of the weight decay on the embedding norms in SimCLR and SimSiam. Epochs are log-scale and go from 1 to 256. Corresponding $k$NN accuracies are plotted in the bottom row.
Cut-Initialization
As shown in Figure 5, weight decay gradually reduces embedding norms over time, but doesn’t control them at the start of training. To address this, we propose cut-initialization — dividing all network weights by a constant $c$ at initialization. This ensures small embedding norms at initialization which are then kept small via weight decay. We implement this uniformly across all layers for simplicity (Listing LABEL:alg:cut_init). Interestingly, a variant of this can be found in HuggingFace’s default image transformer code (PyTorch Foundation, 2025) and we know at least one cos.sim.-based SSL model which uses it (Peng et al., 2022).
@@
We study the interplay between cut-initialization and weight decay values on SimCLR and SimSiam in Table 2. Specifically, we report the $k$NN classification accuracy after 100 epochs on the Cifar-100 and ImageNet-100 datasets and see that, in both the contrastive and non-contrastive settings, pairing cut-initialization with weight decay accelerates the training process. As was the case for the weight decay, the difference is stronger in the non-contrastive setting. $c=2$ and $c=4$ performed best for SimCLR, providing an additional 1-2% in accuracy at the default weight decay, while $c=8$ led to about a 10% increase for SimSiam. Extrapolating from the trends in Table 2, we use $c=3$ for SimCLR and $c=9$ for SimSiam going forward. These are futher analyzed in Figure S4, which also includes BYOL experiments. We show the $k$NN classifier accuracies at $500$ epochs with and without cut-initialization in Table 4. Here we see that pairing SSL models with cut-initialization often helps the model reach higher final accuracies.
We also evaluate cut-initialization in imbalanced data settings. For Cifar-10, we use the exponential split from Van Assel & Balestriero (2024), where class $i$ has $n_{i}=5000\cdot 1.5^{-i}$ samples. Similarly for Cifar-100, the $i$-th class receives $n_{i}=500\cdot 1.034^{-i}$ samples. This way, all classes are represented and both imbalanced datasets contain roughly 15K samples. We also use Flowers102’s naturally long-tailed test set for training (Nilsback & Zisserman, 2008), evaluating on its validation set. Table 5 then shows that, in class-imbalanced settings, pairing SSL training with interventions on the embedding-norm effect can provide double-digit accuracy improvements.
Lastly, Table 3 shows that cut-initialization also provides the expected boost in performance for MoCov2 and MoCov3 (Chen et al., 2020b; Fan et al., 2021), which are cos.sim.-based but use a transformer backbone (Dosovitskiy et al., 2020). We point out that Dino’s objective function does not use the cosine similarity (Caron et al., 2021). Consequently, it does not benefit from cut-initialization.
Table 2: $k$NN accuracy at epoch 100 for various values of cut-constant $c$ and weight decay $\lambda$, color-coded by value. Left: SimCLR on Cifar-100. Right: SimSiam on ImageNet-100. Default weight-decay is underlined.
@@
Table 3: Imagenet-100 $k$NN accuracy (epoch 100) for MoCo/Dino.
Top-1 Top-5
MoCo V2 $c=1$ 38.7 69.7
$c=3$ 43.8 71.8
MoCo V3 $c=1$ 54.8 82.7
$c=3$ 58.8 86.0
Dino $c=1$ 43.7 76.6
$c=3$ 29.0 55.6
GradScale Layer
Perhaps the cleanest way to overcome the embedding-norm effect is to simply rescale the gradient directly. We achieve this by introducing a custom PyTorch autograd.Function which we refer to as GradScale (for a full implementation, see Listing LABEL:alg:grad_scaling in the Appendix). This layer accepts a power parameter $p$ and is simply the identity function in the forward pass. However, the backwards pass multiplies each sample $z_{i}$’s contribution to the gradient by $\|z_{i}\|^{p}$. Thus, choosing power $p=0$ gives the default setting while choosing $p=1$ will cancel the gradients’ dependence on the embedding norm. We visualize the resulting gradient fields with powers $p=0$ and $p=1$ for a 2D embedding space in Figure 6. We refer to training with GradScale power $p=1$ simply as GradScale. This can also be interpreted as a per-point adaptive learning rate that is scaled by each point’s embedding norm.
@@
Table 4 demonstrates that, when training remains stable, SimCLR’s $k$NN accuracy benefits from GradScale’s embedding norm cancellation. Consistent with our cut-initialization experiments, this improvement becomes particularly pronounced on the class-imbalanced datasets in Table 5: GradScale provides a roughly 5% accuracy increase across the imbalanced datasets. While these results are promising, we observed that GradScale with $p=1$ failed to converge on Imagenet-100 and for non-contrastive models. This limitation aligns with these models’ known sensitivity issues (Zhang et al., 2022; Richemond et al., 2023).
Figure 6: Left: the default gradient field of the cosine similarity with respect to the north-pointing line (orange). Right: the gradient field using the GradScale layer with $p=1$.
Table 4: $k$NN accuracies at epoch 1000 for Cifar-10/100 and at epoch 500 for ImageNet/TinyImagenet for default, cut-initialized and GradScale training on standard image datasets.
Cifar 10 Cifar 100 Imagenet 100 Tiny Imagenet
SimCLR Default 87.7 56.6 59.4 36.0
Cut ($c=3$) 88.2 60.1 60.9 37.7
GradScale 88.2 58.2 01.0 38.1
SimSiam Default 88.1 61.8 62.0 41.4
Cut ($c=9$) 88.4 62.6 67.2 41.7
Table 5: $k$NN accuracies at epoch 500 for default, cut-initialized and GradScale training on class-imbalanced image datasets.
@@
Embedding-norm Effect Under Adaptive Optimizers
A potential concern is that our theoretical analysis, which focuses on standard gradient descent, may not extend to adaptive optimizers. To address this, we evaluate our embedding norm mitigation strategies using the Adam optimizer (Kingma, 2014). As shown in Table 6, the embedding-norm effect persists under Adam optimization: both cut-initialization and GradScale continue to provide consistent improvements across SimCLR and SimSiam on both Cifar-10 and Cifar-100. Table S4 reveals that our mitigation strategies have the expected effects on embedding magnitudes even under adaptive optimization: cut-initialization ensures small norms ($\|z\|=2.1$) while GradScale allows them to grow ($\|z\|=175$) without affecting convergence.
Takeaways
Table 6: $k$NN accuracies at epoch 100 using the Adam optimizer for SimCLR and SimSiam under embedding norm interventions.
Cifar-10 Cifar-100
SimCLR Default 79.6 44.3
Cut 79.9 45.3
GradScale 80.5 46.6
SimSiam Default 73.7 36.4
Cut 79.8 44.9
These results make it clear that the embedding norm effect impacts SSL training — particularly in non-contrastive models — and can be mitigated using our proposed strategies. The effect appears most detrimental in class-imbalanced settings, aligning with our results on SSL confidence: imbalanced data creates variance in embedding norms, destabilizing training. Nonetheless, there remain questions which are beyond the scope of this work:
@@
In addition to seeing that the embedding-norm effect is more pronounced in attraction-only settings, we have found that the embeddings can shrink even in the absence of weight decay. We attribute the latter phenomenon to the network’s shared weights: while our theory predicts uniform embedding growth, producing these embeddings via a single set of weights creates competition between different regions of the latent space.
Finally, we have been careful to not describe the embedding-norm effect as a strictly negative phenomenon. Consider the common transfer-learning setting in which prototypical class examples should anchor the learned representation (Pan et al., 2019; Lee et al., 2022). Our findings suggest the embedding-norm effect may naturally support this goal: prototypical examples get large embedding norms and consequently receive smaller gradients.
@@ 7 Discussion
Given that these properties are natively present in widely-used SSL approaches, they suggest several research directions beyond those already stated in the paper. First, embedding norms could serve as simple yet effective reliability metrics during inference. Second, the embedding norms provide a new lens for studying the modality gap in multi-modal representation learning (Liang et al., 2022). Lastly, the interplay between the embedding norm’s roles — as both a confidence metric and gradient scalar — suggests training pipelines that explicitly leverage both mechanisms.
Impact Statement
The goal of our paper is to advance the field of machine learning. We do not see any potential societal consequences of our work that would be worth specifically highlighting.
@@ Acknowledgements
Andrew Draganov is partially supported by the Independent Research Fund Denmark (DFF) under a Sapere Aude Research Leader grant No 1051-00106B, by a StiboFund IT travel grant for PhD students and by project W2/W3-108 Impuls und Vernetzungsfonds der Helmholtz-Gemeinschaft. Sharvaree Vadgama is supported by the Hybrid Intelligence Center, a 10-year program funded by the Dutch Ministry of Education, Culture and Science through the Netherlands Organisation for Scientific Research (NWO). This work was partially funded by the Gemeinnützige Hertie-Stiftung, the Cyber Valley Research Fund (D.30.28739), and the National Institutes of Health (UM1MH130981). The content is solely the responsibility of the authors and does not necessarily represent the official views of the National Institutes of Health. Dmitry Kobak is a member of the Germany’s Excellence cluster 2064 “Machine Learning — New Perspectives for Science” (EXC 390727645). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Jan Niklas Böhm.
@@ A.2 InfoNCE Gradients
$\nabla_{i}^{\mathcal{R}}$ $=-\frac{\sum_{k\neq i}\text{ExpSim}(z_{i},z_{k})\frac{\partial\frac{z_{i}^{\top}z_{k}}{\|z_{i}\|\cdot\|z_{k}\|}}{\partial z_{i}}}{\sum_{k\neq i}\text{ExpSim}(z_{i},z_{k})}$
$=-\frac{\sum_{k\neq i}\text{ExpSim}(z_{i},z_{k})\frac{\partial\frac{z_{i}^{\top}z_{k}}{\|z_{i}\|\cdot\|z_{k}\|}}{\partial z_{i}}}{S_{i}}$
@@
We sum this with the gradients of the attractive term to obtain the full InfoNCE gradient for a positive sample.
Similarly, for $z_{l}$ we obtain
$\nabla_{l}=\nabla_{l}^{\mathcal{R}}=-\frac{\text{ExpSim}(z_{i},z_{l})}{\sum_{k\neq i}\text{ExpSim}(z_{i},z_{k})}\frac{\partial\frac{z_{i}^{\top}z_{l}}{\|z_{i}\|\cdot\|z_{l}\|}}{\partial z_{l}}=-\frac{\text{ExpSim}(z_{i},z_{l})}{S_{i}\|z_{l}\|}(\hat{z}_{i})_{\perp z_{l}},$ (6)
where the last equality follows again from Prop. 1. ∎
We note that the gradients of the InfoNCE loss on positive and negative samples are always tangent to the respective points.
@@
First, consider that we applied the cosine similarity’s gradients from Proposition 1. Since $z_{i}$ and $(z_{j})_{\perp z_{i}}$ are orthogonal, $\|z_{i}^{\prime}\|_{2}^{2}=\|z_{i}\|^{2}+\frac{\gamma^{2}}{\|z_{i}\|^{2}}\|(z_{j})_{\perp z_{i}}\|^{2}$. The second term is positive if $\sin\phi_{ij}>0$, which is the case save for identical or antipodal points, because then, we have $\phi_{ij}\in(0,\pi)$.
@@ B.1 Nonparametric Simulations
Specifically, the mean angle among positive pairs is controlled using the $\alpha$ parameter. In essence, when $\alpha=0$, $x_{2}^{\prime}=x_{2}$. However, when $\alpha=1$ (resp. $\alpha=-1$), $x_{2}^{\prime}=x_{1}$ (resp. $x_{2}^{\prime}=-x_{1}$).
@@ B.2 Parametric Simulations
We restate the entire implementation for the simulations in Section 4.2 for completeness. We choose centers for 4 latent classes $\{c_{1},c_{2},c_{3},c_{4}\}$ uniformly at random from $\mathbb{S}^{10}$ by randomly sampling vectors from a standard multivariate normal distribution and normalizing them to the hypersphere. We then obtain the latent samples $\tilde{z}$ around center $c_{i}$ via $z\sim\mathcal{N}(c_{i},0.1\cdot\mathbf{I})$ and re-normalizing to the hypersphere. For each center, we produce 1000latent samples; these constitute our latent classes. We depict an example of 8 such latent classes (in 3 dimensions) in Figure 2(a). We finally obtain the dataset by generating a random matrix in $\mathbb{R}^{11\times 64}$ and applying it to the latent samples.